OpenStudy (anonymous):

Anyone good with radical equations?? I need your help in working out this problem so I can understand how to do it! n square root 2 = square root 9 - 3n

OpenStudy (jdoe0001):

\(\bf n\sqrt{2}=\sqrt{9}-3n\implies n\sqrt{2}+3n=\sqrt{9}\implies n(\sqrt{2}+3)=\sqrt{9} \\ \quad \\ n=\cfrac{\sqrt{9}}{(\sqrt{2}+3)}\)

OpenStudy (anonymous):

Thanks so much =) but a question for the answer it has 1.5 ? But I get 2 so is it possible the answer 1.5 is wrong?

OpenStudy (jdoe0001):

based on what you gave, the procedure is correct